OpenStudy (yaya090600):

Find the sum of a finite geometric sequence from n = 1 to n = 5, using the expression −3(4)^n − 1. 1,023 1,223 −1,023 −4,374

OpenStudy (dinamix):

\[S _{n} =a _{p} \frac{ 1-r ^{r-p} }{ 1-r }\] use this @yaya090600

OpenStudy (yaya090600):

how do i plug it in?

OpenStudy (dinamix):

\[S_{n} = a_{p}\frac{ 1-r ^{n-p} }{ 1-r}\]

OpenStudy (dinamix):

p=1 n= 5 r=4 a1calculate with use expression -3(4)^n -1

hartnn (hartnn):

not sure why the exponent in the numerator is n-p, instead of just n

hartnn (hartnn):

use this formula \(\Large S_n = a_1 \dfrac{r^n-1}{r-1}\) a1 = 1st term =-3 r = common ratio = 4 n= 5

OpenStudy (dinamix):

but i 'm used \[a_{0}\] first term @hartnn

OpenStudy (dinamix):

this how i get this formula

hartnn (hartnn):

oh! so p= 0 :)

OpenStudy (dinamix):

yup @hartnn

OpenStudy (dinamix):

;D

hartnn (hartnn):

with p=0, our formula are same :)

OpenStudy (dinamix):

@hartnn a0 is first term not a1

hartnn (hartnn):

right, when you start n from 0 and use the formula you posted.

OpenStudy (dinamix):

yup @hartnn

hartnn (hartnn):

a1 is the first term when we start n from 1 and use the formula i posted.

OpenStudy (yaya090600):

I put A, is that correct?

OpenStudy (dinamix):

\[S _{5} = -13 \frac{ 1-4^4 }{ 1-4} =-13\frac{ -255 }{ -3 } =-1105\] the answer will be across @hartnn

OpenStudy (yaya090600):

thats not one of the choices

hartnn (hartnn):

so, using my formula \(S_n = -3 \dfrac{4^5 -1}{4-1} = 1-4^5 = -1023\)

OpenStudy (yaya090600):

so its C

OpenStudy (dinamix):

yup its is @yaya090600 sorry if i was late

OpenStudy (yaya090600):

Thank you guys so much

OpenStudy (dinamix):

@hartnn there sum from n=0 to n=5

OpenStudy (dinamix):

there

OpenStudy (dinamix):

@@yaya090600 u understand or no ?